# The failure of Hölder regularity of solutions for the Camassa–Holm type equation in Besov spaces

### 20 Jan 2024 v1

@@ 3.1 Construction of Initial Data
We need to introduce smooth, radial cut-off functions to localize the frequency region. Precisely, let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
@@
Due to the fact $\varphi(\xi)\equiv 1$ for $\frac{4}{3}\leq|\xi|\leq\frac{3}{2}$, namely,
$\varphi(2^{-j}\xi)\equiv 1\quad\text{in}\;\left\{\xi\in\mathbb{R}^{d}:\ \frac{4}{3}2^{j}\leq|\xi|\leq\frac{3}{2}2^{j}\right\},$
we have
@@
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some sufficiently enough $\delta>0$ such that
$\|u_{0}\|_{B^{s}_{p,r}}\leq C,$
$\|u_{0}^{k}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}\geq cn^{-2}2^{n(1-s)},\quad k\in\{1,2\}$
@@
Proof.  Using (3.8) yields
@@
By the definition of ${B}_{p,r}^{s}$ and , we deduce that
$\|u_{0}\|_{{B}_{p,r}^{s}(\mathbb{R})}$ $=\left\|2^{js}\|\Delta_{j}u_{0}\|_{L^{p}(\mathbb{R})}\right\|_{\ell^{r}(j\geq 1)}\leq\left\|\frac{1}{j^{2}}\right\|_{\ell^{r}(j\geq 1)}\|\phi\|_{L^{p}(\mathbb{R})}\leq C\|\phi\|_{L^{p}(\mathbb{R})}.$
@@
$n^{2}u_{0}^{k}\partial_{x}\Delta_{n}u_{0}$ $=2^{-ns}u_{0}^{k}(x)\phi^{\prime}(x)\cos\left(\frac{17}{12}2^{n}x\right)-\frac{17}{12}2^{n}2^{-ns}u_{0}^{k}(x)\phi(x)\sin\left(\frac{17}{12}2^{n}x\right).$
Since $u_{0}^{k}(x)$ is a real-valued and continuous function on $\mathbb{R}$, then there exists some $\delta>0$ such that
$|u_{0}^{k}(x)|\geq\frac{1}{2}|u_{0}^{k}(0)|=\frac{1}{2}\left(\phi(0)\sum\limits^{\infty}_{n=3}n^{-2}2^{-ns}\right)^{k}=:c_{0}\quad\text{ for any }x\in B_{\delta}(0).$ (3.11)
@@
$n^{2}\|u_{0}^{k}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}$ $\geq c_{0}2^{n}2^{-ns}\left\|\phi(x)\sin\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(B_{\delta}(0))}-C2^{-ns}\left\|\phi^{\prime}(x)\phi^{k}(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}$
@@
We choose $n$ large enough such that $C<\frac{c}{2}2^{n}$ and then finish the proof of Lemma 3.1.
@@ 3.2 Error Estimates
$\|\mathbf{S}_{t}\left(u_{0}\right)-u_{0}\|_{B^{s-1}_{p,r}}\lesssim t.$ (3.12)
@@
    $\|\mathbf{w}\|_{B^{s-2}_{p,r}}\lesssim t^{2};$
    here and in what follows we denote
[…]
  • $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$, we have
@@
Proof.  For simplicity, we denote $u(t):=\mathbf{S}_{t}(u_{0})$ here and in what follows. Notice that $(s,p,r)$ satisfies (1.4), then using the local well-posedness result, we know that there exists a positive time $T$ such that $u(t)\in\mathcal{C}([0,T];B_{p,r}^{s})$. Furthermore, it holds that
$\|u(t)\|_{L^{\infty}_{T}B^{s}_{p,r}}\leq C\|u_{0}\|_{B^{s}_{p,r}}\leq C.$
Noticing that $u_{0}\in B^{s}_{\infty,r}$, we can deduce that $u(t)\in\mathcal{C}([0,T];B_{\infty,r}^{s})$ and also hold
$\|u(t)\|_{L^{\infty}_{T}B^{s}_{\infty,r}}\leq C\|u_{0}\|_{B^{s}_{\infty,r}}\leq C.$
Using the Newton-Leibniz formula and Remark 2.1 with $v=0$, we obtain from (1.2) that
$\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\leq\int^{t}_{0}\|\mathbf{P}(u)\|_{B^{s-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim t\|u\|^{2}_{L_{t}^{\infty}B^{s}_{p,r}}$
$\lesssim t\|u_{0}\|^{2}_{B^{s}_{p,r}}$
t.less-than-or-similar-$\lesssim t.$ (3.13)
@@
$\|\mathbf{w}\|_{B^{s-2}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u-\mathbf{v}_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{s-2}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\|u+u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim t^{2}.$ (3.14)
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$.
Then using the embedding $B^{s-2}_{p,r}\hookrightarrow B^{-1}_{p,r}$, we have
$\|\mathbf{w}\|_{B^{0}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u-\mathbf{v}_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{0}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{-1}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|(u-u_{0})(u+u_{0})\|_{B^{1}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\underbrace{\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}}_{=:\ I_{1}}\mathrm{d}\tau+\int^{t}_{0}\underbrace{\|u-u_{0}\|_{B^{1}_{p,r}}\|u+u_{0}\|_{B^{1}_{p,r}}}_{=:\ I_{2}}\mathrm{d}\tau.$ (3.15)
@@
$\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})=T_{\partial_{x}(u-u_{0})}\partial_{x}(u+u_{0})+T_{\partial_{x}(u+u_{0})}\partial_{x}(u-u_{0})+R\big{(}\partial_{x}(u+u_{0}),\partial_{x}(u-u_{0})\big{)}.$
@@
$\|T_{\partial_{x}(u-u_{0})}\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u-u_{0})\|_{B^{-1}_{\infty,\infty}}\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{p,r}}$
$\lesssim\|u-u_{0}\|_{B^{1}_{p,r}}\|u+u_{0}\|_{B^{s}_{p,r}}$
$\lesssim t^{s-1},$
$\|T_{\partial_{x}(u+u_{0})}\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u+u_{0})\|_{L^{\infty}}\|\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}$
$\lesssim\|u+u_{0}\|_{B^{s-1}_{p,r}}\|u-u_{0}\|_{B^{s-1}_{p,r}}$
t,less-than-or-similar-$\lesssim t,$
$\|R\big{(}\partial_{x}(u+u_{0}),\partial_{x}(u-u_{0})\big{)}\|_{B^{s-1}_{p,r}}$ $\lesssim\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{\infty,\infty}}\|\partial_{x}(u-u_{0})\|_{B^{0}_{p,r}}$
$\lesssim\|u+u_{0}\|_{B^{s}_{\infty,r}}\|u-u_{0}\|_{B^{1}_{p,r}}$
$\lesssim t^{s-1},$
where we have used the fact
$\|u(t)\|_{B^{s}_{\infty,r}}\lesssim\|u_{0}\|_{B^{s}_{\infty,r}}\lesssim 1$
and the interpolation argument
$\|u-u_{0}\|_{B^{1}_{p,r}}\lesssim\|u(t)-u_{0}\|^{s-1}_{B^{s-1}_{p,r}}\|u(t)-u_{0}\|^{2-s}_{B^{s}_{p,r}}\lesssim t^{s-1}.$
@@
$I_{2}\lesssim\|u-u_{0}\|_{B^{1}_{p,r}}\lesssim t^{s-1}.$ (3.17)
Inserting (3.16) and (3.17) into (3.2) yields the desired result. Thus, we finish the proof of Proposition 3.1.
3.3 Proof of Theorem 1.1
Now we present the proof of Theorem 1.1. Notice that
$\mathbf{S}_{t}(u_{0})-u_{0}=t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\quad\text{and}\quad\mathbf{\widetilde{u}}_{0}=\mathbf{P}(u_{0})-u_{0}\partial_{x}u_{0}.$
Case 1: $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty)$.
By the triangle inequality and Propositions 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(u_{0})-u_{0}\big{)}\big{\|}_{L^{p}}$
$=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\big{)}\big{\|}_{L^{p}}$
@@
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}$
$\quad-Ct\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}.$ (3.18)
By Lemmas 2.2-2.3, one has
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}\lesssim\|(\partial_{x}u_{0})^{2}+u_{0}^{2}\|_{B^{s-1}_{p,r}}\lesssim 1$
and by Lemma 2.5
@@
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{n}-Ct^{1-\alpha}-C2^{2{n}}t^{2-\alpha},$
@@
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$.
By the triangle inequality and Propositions 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big{\|}\Delta_{n}\big{(}\mathbf{S}_{t}(u_{0})-u_{0}\big{)}\big{\|}_{L^{p}}$
$=2^{{ns}}\big{\|}\Delta_{n}\big{(}t\mathbf{v}_{0}+\mathbf{w}\big{)}\big{\|}_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{v}_{0}\|_{L^{p}}-2^{{ns}}\|\mathbf{w}\|_{B^{0}_{p,r}}$
$\geq t2^{{n}s}\|\Delta_{n}\big{(}u_{0}\partial_{x}u_{0}\big{)}\|_{L^{p}}-t2^{{n}s}\|\Delta_{n}\big{(}\mathbf{P}(u_{0})\big{)}\|_{L^{p}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big{\|}2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big{\|}_{\ell^{\infty}}$
$\quad-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C(2^{{n}}t)^{s},$ (3.19)
@@
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}\lesssim\|(\partial_{x}u_{0})^{2}+u_{0}^{2}\|_{B^{s}_{p,r}}\lesssim 1$
and
@@
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{{n}}-Ct^{1-\alpha}-C(2^{{n}}t)^{s}t^{-\alpha},$

### 14 Mar 2026 v2

@@ 3.1 Construction of Initial Data
Let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
@@
Due to the fact $\varphi(\xi)\equiv 1$ for $\frac{4}{3}\leq|\xi|\leq\frac{3}{2}$, we have
@@
Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ such that
$\|u_{0}\|_{B^{s}_{p,r}}\leq C\quad\text{and}\quad\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}\geq cn^{-2}2^{n(1-s)},$
@@
Proof.
Using (3.4) yields
@@
By the definition of ${B}_{p,r}^{s}$, we deduce that
$\|u_{0}\|_{{B}_{p,r}^{s}(\mathbb{R})}$ $=\left\|2^{js}\|\Delta_{j}u_{0}\|_{L^{p}(\mathbb{R})}\right\|_{\ell^{r}(j\geq 1)}\leq\left\|j^{-2}\right\|_{\ell^{r}(j\geq 1)}\|\phi\|_{L^{p}(\mathbb{R})}\leq C\|\phi\|_{L^{p}(\mathbb{R})}.$
@@
$n^{2}u_{0}\partial_{x}\Delta_{n}u_{0}$ $=2^{-ns}u_{0}(x)\phi^{\prime}(x)\cos\left(\frac{17}{12}2^{n}x\right)-\frac{17}{12}2^{n}2^{-ns}u_{0}(x)\phi(x)\sin\left(\frac{17}{12}2^{n}x\right).$
Since $u_{0}(x)$ is a real-valued and continuous function on $\mathbb{R}$, then there exists some $\delta>0$ such that for any $x\in B_{\delta}(0)$
$|u_{0}(x)|\geq\frac{1}{2}|u_{0}(0)|=\frac{1}{2}\phi(0)\sum\limits^{\infty}_{n=3}n^{-2}2^{-ns}=:c_{0}.$ (3.7)
@@
$n^{2}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}$ $\geq c_{0}2^{n}2^{-ns}\left\|\phi(x)\sin\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(B_{\delta}(0))}-C2^{-ns}\left\|\phi^{\prime}(x)\phi(x)\cos\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}}$
@@
We choose $n$ large enough such that $C<\frac{c}{2}2^{n}$ and then finish the proof of Lemma 3.1. ∎
@@ 3.2 Error Estimates
$\|\mathbf{S}_{t}\left(u_{0}\right)-u_{0}\|_{B^{s-1}_{p,r}}\lesssim t.$
@@
    $\|\mathbf{w}\|_{B^{s-2}_{p,r}}\lesssim t^{2},$
    here and in what follows we denote $\mathbf{w}:=\mathbf{S}_{t}(u_{0})-u_{0}-t\mathbf{\widetilde{u}}_{0}$ with $\mathbf{\widetilde{u}}_{0}:=\mathbf{P}(u_{0})-u_{0}\partial_{x}u_{0}.$
  • for $s\in\big(1+\frac{1}{p},\frac{3}{2}\big]$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$, we have
@@
Proof.
For simplicity, we denote $u(t):=\mathbf{S}_{t}(u_{0})$ here and in what follows. Notice that $(s,p,r)$ satisfies (1.3) and by the local well-posedness result [7, 13, 16], there exists a positive time $T$ such that $u(t)\in\mathcal{C}([0,T];B_{p,r}^{s})$. Furthermore, it holds that $\|u(t)\|_{L^{\infty}_{T}B^{s}_{p,r}}\leq C\|u_{0}\|_{B^{s}_{p,r}}\leq C.$ For any $u_{0}\in B^{s}_{\infty,r}$, we can deduce that $u(t)\in\mathcal{C}([0,T];B_{\infty,r}^{s})$ and $\|u(t)\|_{L^{\infty}_{T}B^{s}_{\infty,r}}\leq C\|u_{0}\|_{B^{s}_{\infty,r}}\leq C.$ Using the Newton-Leibniz formula and Remark 2.1 with $v=0$, we obtain from (1.2) that
$\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}$ $\leq\int^{t}_{0}\|\partial_{\tau}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\leq\int^{t}_{0}\|\mathbf{P}(u)\|_{B^{s-1}_{p,r}}+\|u\partial_{x}u\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t\|u\|^{2}_{L_{t}^{\infty}B^{s}_{p,r}}\lesssim t\|u_{0}\|^{2}_{B^{s}_{p,r}}\lesssim t.$ (3.8)
@@
$\|\mathbf{w}\|_{B^{s-2}_{p,r}}\leq$ $~\int^{t}_{0}\|\partial_{\tau}u-\mathbf{\widetilde{u}}_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{s-2}_{p,r}}+\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{s-2}_{p,r}}+\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|\partial_{x}(u-u_{0})\|_{B^{s-2}_{p,r}}\|\partial_{x}(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\|u+u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$. Using the embedding $B^{s-2}_{p,r}\hookrightarrow B^{-1}_{p,r}$, we have
$\|\mathbf{w}\|_{B^{0}_{p,r}}\leq$ $~\int^{t}_{0}\|\partial_{\tau}u-\mathbf{\widetilde{u}}_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau\lesssim\int^{t}_{0}\|\mathbf{P}(u)-\mathbf{P}(u_{0})\|_{B^{0}_{p,r}}+\|u\partial_{x}u-u_{0}\partial_{x}u_{0}\|_{B^{0}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|\partial_{x}(u-u_{0})\partial_{x}(u+u_{0})\|_{B^{-1}_{p,r}}+\|(u-u_{0})(u+u_{0})\|_{B^{1}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\underbrace{\|\partial_{x}(u(\tau)-u_{0})\partial_{x}(u(\tau)+u_{0})\|_{B^{s-2}_{p,r}}}_{=:\ I_{1}(\tau)}+\underbrace{\|u(\tau)-u_{0}\|_{B^{1}_{p,r}}\|u(\tau)+u_{0}\|_{B^{1}_{p,r}}}_{=:\ I_{2}(\tau)}\mathrm{d}\tau.$ (3.9)
@@
$\partial_{x}(u(t)-u_{0})\partial_{x}(u(t)+u_{0})$ $=T_{\partial_{x}(u(t)-u_{0})}\partial_{x}(u(t)+u_{0})+T_{\partial_{x}(u(t)+u_{0})}\partial_{x}(u(t)-u_{0})+R\big(\partial_{x}(u(t)+u_{0}),\partial_{x}(u(t)-u_{0})\big).$
@@
$\|T_{\partial_{x}(u(t)-u_{0})}\partial_{x}(u(t)+u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u(t)-u_{0})\|_{B^{-1}_{\infty,\infty}}\|\partial_{x}(u(t)+u_{0})\|_{B^{s-1}_{p,r}}$
$\lesssim\|u(t)-u_{0}\|_{B^{1}_{p,r}}\|u(t)+u_{0}\|_{B^{s}_{p,r}}\lesssim t^{s-1},$
$\|T_{\partial_{x}(u(t)+u_{0})}\partial_{x}(u(t)-u_{0})\|_{B^{s-2}_{p,r}}$ $\lesssim\|\partial_{x}(u(t)+u_{0})\|_{L^{\infty}}\|\partial_{x}(u(t)-u_{0})\|_{B^{s-2}_{p,r}}$
$\lesssim\|u(t)+u_{0}\|_{B^{s-1}_{p,r}}\|u(t)-u_{0}\|_{B^{s-1}_{p,r}}\lesssim t,$
$\|R\big(\partial_{x}(u(t)+u_{0}),\partial_{x}(u(t)-u_{0})\big)\|_{B^{s-1}_{p,r}}$ $\lesssim\|\partial_{x}(u(t)+u_{0})\|_{B^{s-1}_{\infty,\infty}}\|\partial_{x}(u(t)-u_{0})\|_{B^{0}_{p,r}}$
$\lesssim\|u(t)+u_{0}\|_{B^{s}_{\infty,r}}\|u(t)-u_{0}\|_{B^{1}_{p,r}}\lesssim t^{s-1},$
where we have used the fact $\|u(t)\|_{B^{s}_{\infty,r}}\lesssim\|u_{0}\|_{B^{s}_{\infty,r}}\lesssim 1$ and the interpolation argument
$\|u(t)-u_{0}\|_{B^{1}_{p,r}}\lesssim\|u(t)-u_{0}\|^{s-1}_{B^{s-1}_{p,r}}\|u(t)-u_{0}\|^{2-s}_{B^{s}_{p,r}}\lesssim t^{s-1}.$
@@
$I_{2}(t)\lesssim\|u(t)-u_{0}\|_{B^{1}_{p,r}}\lesssim t^{s-1}.$ (3.11)
Inserting (3.10) and (3.11) into (3.2) yields the desired result. Thus, we finish the proof of Proposition 3.1. ∎
Now we present the proof of Theorem 1.1.
Proof.
Case 1: $s>\max\left\{1+\frac{1}{p},\frac{3}{2}\right\}$ and $(p,r)\in[1,\infty]\times[1,\infty)$. Notice that $\mathbf{S}_{t}(u_{0})-u_{0}=t\mathbf{\widetilde{u}}_{0}+\mathbf{w}$ and $\mathbf{\widetilde{u}}_{0}=\mathbf{P}(u_{0})-u_{0}\partial_{x}u_{0}.$ By the triangle inequality and Proposition 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big\|\Delta_{n}\big(\mathbf{S}_{t}(u_{0})-u_{0}\big)\big\|_{L^{p}}=2^{{ns}}\big\|\Delta_{n}\big(t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\big)\big\|_{L^{p}}$
@@
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}-Ct\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big\|2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big\|_{\ell^{\infty}}-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C2^{2{n}}t^{2}.$ (3.12)
By Lemmas 2.2-2.4, one has
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}\lesssim\left\|u_{0}^{2}+\frac{1}{2}(\partial_{x}u_{0})^{2}\right\|_{B^{s-1}_{p,r}}\lesssim 1,$
@@
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{n}-Ct^{1-\alpha}-C2^{2{n}}t^{2-\alpha}.$
@@
Case 2: $1+\frac{1}{p}<s\leq\frac{3}{2}$ and $(p,r)\in[1,\infty]\times[1,\infty)$ or $s=1+\frac{1}{p}$ and $(p,r)\in[1,\infty)\times\{1\}$. By Proposition 3.1, we deduce that
$\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}$ $\geq 2^{{ns}}\big\|\Delta_{n}\big(\mathbf{S}_{t}(u_{0})-u_{0}\big)\big\|_{L^{p}}=2^{{ns}}\big\|\Delta_{n}\big(t\mathbf{\widetilde{u}}_{0}+\mathbf{w}\big)\big\|_{L^{p}}$
$\geq t2^{{ns}}\|\Delta_{n}\mathbf{\widetilde{u}}_{0}\|_{L^{p}}-2^{{ns}}\|\mathbf{w}\|_{B^{0}_{p,r}}\geq t2^{{n}s}\|\Delta_{n}\big(u_{0}\partial_{x}u_{0}\big)\|_{L^{p}}-t2^{{n}s}\|\Delta_{n}\big(\mathbf{P}(u_{0})\big)\|_{L^{p}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-t2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}-C(2^{{n}}t)^{s}$
$\geq t2^{{n}s}\|u_{0}\partial_{x}\Delta_{n}u_{0}\|_{L^{p}}-Ct\big\|2^{{n}s}\|[\Delta_{n},u_{0}]\partial_{x}u_{0}\|_{L^{p}}\big\|_{\ell^{\infty}}-t\|\mathbf{P}(u_{0})\|_{B^{s}_{p,\infty}}-C(2^{{n}}t)^{s},$ (3.13)
@@
$\|\mathbf{P}(u_{0})\|_{B^{s}_{p,r}}\lesssim\left\|u_{0}^{2}+\frac{1}{2}(\partial_{x}u_{0})^{2}\right\|_{B^{s-1}_{p,r}}\lesssim 1,$
@@
$t^{-\alpha}\|\mathbf{S}_{t}(u_{0})-u_{0}\|_{B^{s}_{p,r}}\geq ct^{1-\alpha}n^{-2}2^{{n}}-Ct^{1-\alpha}-C(2^{{n}}t)^{s}t^{-\alpha}.$